If a vat is 15% full of oil, and adding an additional 144 liters of oil makes it 75% full, then how much oil is in the container when it is full?

Answer by checkley79(3341)

.15X+144=.75X
.15X-.75X=-144
-.60X=-144
X=-144/-.60
X=240 LITERS IS THE CAPACITY OF THE VAT.
PROOF:
.15*240+144=.75*240
36+144=180
180=180